# ReCo-Diff: Explore Retinex-Based Condition Strategy in Diffusion Model for Low-Light Image Enhancement

### v1

@@ 3.2 Retinex-Based Conditional Diffusion Model
$I=\hat{R}\odot L,\quad\hat{I}=\hat{R}\odot\hat{L},\vspace{-0.2cm}$ (5)
where $\odot$ denotes the element-wise multiplication. According to multi-scale Retinex [17], $\hat{R}$ is the Retinex output by removing the lighting effects in $I$ as follows:
$\begin{split}\hat{R}=\log(&I)-\log(L)=\log(I)-\log(\mathcal{G}(I)),\\ \hat{I}&=\mathcal{T}(\exp(\hat{R})),\quad\hat{I}\in(0,1),\end{split}$ (6)
@@
$\hat{I}=\mathcal{F}(I,\hat{R}),\vspace{-0.18cm}$ (7)
where $\mathcal{F}(\cdot)$ denotes the deep network and $\hat{R}$ acts as an auxiliary guidance. Inspired by the improved Retinex model, we consider the enhancement process from two stages: Retinex-based guidance adjustment and conditional image enhancement. Recalling the conditional DDPM model discussed in Sec. 3.1, the enhancement task can be performed due to the superiority of the generative model conditioned by low-light images [34, 14]. However, the noise and artifacts in low-light images inevitably mislead the diffusion model and cause detail loss and unsatisfactory color.
@@
$\hat{I}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(I,\mathcal{F_{A}}(R,L)),\mathcal{F_{A}}(R,L)),\vspace{-0.18cm}$ (8)
where $\bm{\epsilon}_{\theta}(\cdot)$, $\mathcal{F_{R}}(\cdot)$, and $\mathcal{F_{A}}(\cdot)$ denote UNet, RNet, and ANet in Fig. 2, respectively. Combining the promising texture generation capability of the diffusion model and the conscious representation with vivid color and detail of Retinex theory, we develop a Retinex-based condition strategy for the conditional DDPM and formulate our ReCo-Diff. Eq. 8 represents the denoising and refinement process in one iteration, which can be described as two stages, Retinex-based condition learning and Retinex-conditioned refinement. In the first stage, we obtain the decomposition results from a pre-trained DNet and input them into the following ANet to produce Retienx-based conditions, which will be described in Sec. 3.2.1. Then, the conditions will be introduced into the UNet and RNet to achieve conditional image denoising and refinement as elaborated in Sec. 3.2.2.
3.2.1 Retinex-based Condition Learning
As described in Sec. 3.2, the basis of conditional image refinement is to learn high-quality conditions. Inspired by Retinex-based methods [39, 52], we propose the DNet to serve as a pre-processing module and produce initial Retinex-based conditions. Compared to calculating an invariant color map as conditional input [34, 49], the decomposed maps contain not only color information but also details and illumination, achieving a more sufficient form of conditional inputs. To further exploit the progressively refined images in the diffusion process, DNet inputs $\mathbf{x}_{t}$ as well and provides competent updated predictions $R_{t-1}$ and $L_{t-1}$ for better generation. However, the images in the early steps contain incorrect contents and mislead the decomposition. Therefore, an adjustment network (ANet) is introduced to deal with corruption and ameliorate the low-quality conditions in the early steps, achieving approving conditional enhancement during the whole generative process. Thus, we realize the Retinex-based condition learning as:
$\hat{\mathbf{c}}_{t-1}=\mathcal{F_{A}}(\mathbf{c}_{t-1})=\mathcal{F_{A}}(\mathcal{F_{D}}(\mathbf{y},\mathbf{x}_{t})).$ (9)
Pre-trained DNet. DNet adopts a lightweight UNet-like network to learn the decomposition mapping. It inputs the images $I$ and outputs the reflectance $R$ and illumination $L$. Inspired by the training strategy in [39], we utilize the constant reflectance loss, smooth illumination loss, and reconstruction loss to pre-train DNet. More details can be seen in the supplementary materials.
@@ 3.2.1 Retinex-based Condition Learning
$\mathcal{L}_{ja}=\parallel{R_{t}}\!^{\prime}-\hat{R}\parallel_{1}+SSIM({R_{t}}\!^{\prime},\hat{R}).$ (10)
@@
$\mathcal{L}_{je}=\parallel{L_{t}}\!^{\prime}-\hat{L}\parallel_{1}+Hist({L_{t}}\!^{\prime},\hat{L}),$ (11)
@@
$\mathcal{L}_{ANet}=\lambda_{ja}\mathcal{L}_{ja}+\lambda_{je}\mathcal{L}_{je},$ (12)
@@
3.2.2 Retinex-conditioned Refinement
After obtaining the Retinex-based conditions $\hat{\mathbf{c}}_{t}$, the next stage is to incorporate the conditions into the iterative diffusion process. Previous methods propose to introduce the invariant color maps [49, 34, 57] and Retinex decomposition [48]. However, existing condition strategies are insufficient to conduct the favorable iterative enhancement. In our ReCo-Diff, we already learn the Retinex-based conditions as described in Sec. 3.2.1. Then, following the usage of color map [49, 34], we first control the generation capability of UNet by leveraging $\hat{\mathbf{c}}_{t}$ as extra inputs, while the carefully prepared conditions are being underutilized just for denoising. Thus, as illustrated in Eq. 8, we apply a post-processing refinement network (RNet) to fully exploit the conditions and guarantee the consistent color and contents of ${\hat{\mathbf{x}}_{0}}^{\prime}$. Different from [48], RNet adopts a lightweight architecture and implements Retinex-conditioned refinement at both feature- and image-levels by FRCM and IRCM. Overall, the refinement stage can be elaborated as:
${\hat{\mathbf{x}}_{0}}^{\prime}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(\mathbf{x}_{t-1},\hat{\mathbf{c}}_{t}),\hat{\mathbf{c}}_{t}),$ (13)
@@ 3.2.2 Retinex-conditioned Refinement
Conditional Denoising. UNet introduces reflectance and illumination maps through embedding layers and integrates them in feature space instead of concatenation for effective information interaction. Specifically, we first transform the original ${R_{t}}\!^{\prime}\in\mathbb{R}^{W\times H\times 3}$ and ${L_{t}}\!^{\prime}\in\mathbb{R}^{W\times H\times 1}$ into embedding features $R_{em}$ and $L_{em}$. To maintain the content and color information, the reflectance embedding $R_{em}\in\mathbb{R}^{W\times H\times C}$ keeps the original resolution. Moreover, the illumination embedding is downsampled as $L_{em}\in\mathbb{R}^{W/2\times H/2\times C}$ for computational efficiency thanks to the smooth distribution of lightness. Then, the $R_{em}$ and $L_{em}$ are input to the conditional layers with original and half sizes respectively, guiding the denoising process. Thus, we can approximate $\hat{\mathbf{x}}_{0}$ by updating Eq. 2 as:
$\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\hat{\mathbf{c}}_{t},\gamma_{t})\right).\vspace{-0.15cm}$ (14)
Figure 3: The architecture of the feature/image-level Retinex-conditioned modules (F/IRCM). FRCM inputs multi-scale features $F_{t}$ from ANet and obtains the optimized image feature $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$. Then, IRCM inputs the ${R_{t}}\!^{\prime}$, ${L_{t}}\!^{\prime}$ and the $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to refine the approximated $\hat{\mathbf{x}}_{0}$ a produce the final output ${\hat{\mathbf{x}}_{0}}^{\prime}$.
Conditional Refinement. As shown in Fig. 2, RNet consists of two Retinex-conditioned modules to realize the color recovery and detail enhancement of the approximated $\hat{\mathbf{x}}_{0}$. Specifically, we first implement the Retinex-conditioned refinement in feature space, since $F_{t}$ incorporates rich information of adjustment mapping as illustrated in Sec. 3.2.1. Thus, as depicted in Fig. 3, we unify the sizes of multi-scale features in $F_{t}$ and model them as affine transformation parameters $F_{\gamma}$ and $F_{\beta}$. Then, the feature fusion is carried out by scaling and shifting operations $\mathcal{F}_{T}(\cdot|\cdot)$ as:
$F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}=\mathcal{F}_{T}(F_{\hat{\mathbf{x}}_{0}}|F_{t})=F_{\gamma}\odot W(F_{\hat{\mathbf{x}}_{0}})+F_{\beta},\vspace{-0.15cm}$ (15)
@@
$\begin{split}\Delta\hat{\mathbf{x}}_{0}&=\mathcal{F}_{T}({R_{t}}\!^{\prime}|\mathcal{F}_{T}({L_{t}}\!^{\prime}|\hat{\mathbf{x}}_{0})),\\ {\hat{\mathbf{x}}_{0}}^{\prime}&=\mathcal{F}_{T}(\hat{\mathbf{x}}_{0}|\Delta\hat{\mathbf{x}}_{0})+W(F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}),\end{split}$ (16)
@@
where ${\hat{\mathbf{x}}_{0}}^{\prime}$ is the final refined output of RNet. In RNet, FRCM and IRCM fully exploit the Retinex-based conditions and produce results with consistent color and details. Accordingly, in addition to Eq. 4, we also utilize the constraint of ${\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to optimize the ReCo-Diff as follows:
@@
$\mathcal{L}_{RNet}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\mathbf{x}_{0}-\mathcal{F}_{R}(\hat{\mathbf{x}}_{0},\hat{\mathbf{c}}_{t},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.15cm}$ (17)
4 Experiments
4.1 Experimental Settings
@@
Table 1: Quantitative comparison on the LOL [39] and LOL-v2-real [47] datasets. $\uparrow$ ($\downarrow$) denotes that, larger (smaller) values suggest better quality. The best results are highlighted in bold and the second best results are in underline (Special fonts are also used in Tabs. 2 and 3).
Methods LOL LOL-v2-real
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
LIME [11] TIP’16 16.760 0.560 0.350 117.892 15.240 0.470 0.428 118.171
RetinexNet [39] BMVC’18 16.770 0.462 0.474 113.699 18.371 0.723 0.365 133.905
KinD [52] MM’19 20.870 0.799 0.207 104.632 17.544 0.669 0.375 137.346
DRBN [46] CVPR’20 19.860 0.834 0.155 75.359 20.130 0.830 0.147 60.631
Zero-DCE [9] CVPR’20 14.861 0.562 0.335 101.237 18.059 0.580 0.313 91.939
MIRNet [50] PAMI’22 24.140 0.842 0.131 69.179 20.357 0.782 0.317 49.108
SNR [43] CVPR’22 24.608 0.840 0.151 55.121 21.479 0.848 0.157 54.532
LLFlow [36] AAAI’22 24.999 0.870 0.117 65.143 26.200 0.888 0.137 69.793
PairLIE [7] CVPR’23 19.510 0.736 0.248 100.715 20.357 0.782 0.317 96.911
SMG [44] CVPR’23 23.684 0.826 0.118 58.846 24.620 0.867 0.148 78.582
FourLLIE [24] MM’23 20.222 0.766 0.250 91.793 22.340 0.847 0.051 89.334
Retinexformer [1] ICCV’23 25.153 0.843 0.131 71.148 22.794 0.839 0.171 62.439
Diff-Retinex [48] ICCV’23 21.981 0.863 0.048 47.851 - - - -
DiffLL [14] SIGGRAPH Asia’23 26.336 0.845 0.217 48.114 28.857 0.876 0.207 45.359
PyDiff [57] IJCAI’23 27.088 0.875 0.111 56.215 27.236 0.869 0.211 57.832
ReCo-Diff (Ours) 27.626 0.884 0.090 42.990 29.306 0.906 0.105 46.677
Implementation Details. Our method is trained using an NVIDIA RTX A100 GPU and the training iteration is set to 600k. We use the Adam optimizer with the momentum as $(0.9,0.999)$. The initial learning rate is set to $1\!\times\!10^{-4}$ and decays by a factor of 0.5 after every $1\!\times\!10^{5}$ iterations. The batch size and patch size are set to 8 and 256$\times$256, respectively. We followed the similar architecture of [32] as our denoising U-Net $\bm{\epsilon}_{\theta}$. We used 2000 diffusion steps T for training. In the inference stage, we follow the DDIM [33] and set 8 steps for enhancement.
Datasets. We evaluate the proposed method on several datasets from various scenes, including LOL [39], LOL-v2 [47], and UHD-LL [24]. The LOL dataset [39] includes 485 low/normal-light image pairs for training and 15 pairs for testing. The LOL-v2 dataset [47] can be divided into two parts, the real part is an extension of LOL with 689/100 pairs and the synthetic part includes 900/100 pairs for training and testing. The UHD-LL dataset [24] includes 2000 pairs for training and 150 pairs for testing.
Metrics. To conduct comparison, we employ two full-reference metrics peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) [38], and two perceptual metrics learned perceptual image patch similarity (LPIPS) [51], Fréchet Inception Distance (FID) [12] for evaluation.
Compared Methods. We compare with a rich collection of state-of-the-art methods, including LIME [11], RetinexNet [39], KinD [52], DRBN [46], Zero-DCE [9], EnlightGAN [15], MIRNet [50], SNR [43], LLFlow [36], PairLIE [7], SMG [44], FourLLIE [24], Retinexformer [1], UHDFour [24] and diffusion-based methods including Diff-Retinex [48], DiffLL [14], and PyDiff [57].
4.2 Quantitative Evaluation
Table 2: Quantitative comparison on the LOL-v2-synthetic [47].
Methods LOL-v2-synthetic
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
LIME [11] TIP’16 16.880 0.758 0.104 -
RetinexNet [39] BMVC’18 16.551 0.652 0.379 98.843
KinD [52] MM’19 18.956 0.801 0.262 89.156
DRBN [46] CVPR’20 21.687 0.825 0.174 52.972
Zero-DCE [9] CVPR’20 17.756 0.814 0.168 49.239
MIRNet [50] PAMI’22 21.941 0.876 0.112 38.775
SNR [43] CVPR’22 24.130 0.927 0.032 23.971
LLFlow [36] AAAI’22 23.692 0.911 0.103 32.576
PairLIE [7] CVPR’23 19.074 0.794 0.230 85.209
SMG [44] CVPR’23 25.618 0.905 0.053 23.210
FourLLIE [24] MM’23 24.649 0.919 0.039 26.351
RetinexFormer [1] ICCV’23 25.670 0.928 0.059 22.781
DiffLL [14] SIGGRAPH Asia’23 25.456 0.896 0.102 43.670
PyDiff [57] IJCAI’23 26.038 0.922 0.072 26.854
ReCo-Diff (Ours) 26.701 0.935 0.057 22.755
Figure 4: Visual comparison of our ReCo-Diff and the compared LLIE methods on the LOL dataset.
Figure 5: Visual comparison of our ReCo-Diff and the compared LLIE methods on the UHD-LL dataset.
Tabs. 1, 2 and 3 show the comparisons on LOL, LOL-v2, and UHD-LL. It is clear that our ReCo-Diff achieves consistent and significant performance gain over all competing methods. Specifically, our method provides significant improvement of 0.638dB/0.449dB on LOL/LOL-v2 datasets respectively, establishing the new state-of-the-art with PSNR values of 27.626dB/29.306dB. Furthermore, our method achieves similar performance in terms of SSIM, yielding the best values of 0.884/0.906. On the LOL-v2-synthetic test set, our ReCo-Diff improves the PSNR and SSIM by 0.663 and 0.013 than the second-best method. On UHD-LL, we also achieve considerable performance on PSNR and SSIM of 1.121dB and 0.012, proving our capability of generalization. Although our ReCo-Diff obtains several second-best LPIPS and FID values, it outperforms other diffusion-based LLIE methods in most cases and achieves the best FID on LOL, which indicates the overall performance on perceptual metrics is still competitive. Consequently, the considerable performance shows the capability of suppressing noise and preserving color and details by Retinex-based priors, which demonstrates the effectiveness of our proposed Retinex-based condition strategy in resolving LLIE task.
4.3 Qualitative Evaluation
@@
Table 3: Quantitative comparison on the UHD-LL [24]. To conserve space, we select recent methods in 2023 while ensuring that the results are sufficient to demonstrate our superiority.

### v2

@@ 3.2. Retinex Prior Incorporation
(5) $I=\hat{R}\odot L,\quad\hat{I}=\hat{R}\odot\hat{L},\vspace{-0.15cm}$
where $\odot$ denotes the element-wise multiplication. According to multi-scale Retinex (Jobson et al., 1997), $\hat{R}$ has a different form, which is the Retinex output by removing the lighting effects in $I$ as follows:
(6) $\begin{split}\hat{R}=\log(&I)-\log(L)=\log(I)-\log(\mathcal{G}(I)),\\ \hat{I}&=\mathcal{T}(\exp(\hat{R})),\quad\hat{I}\in(0,1),\vspace{-0.15cm}\end{split}$
@@
(7) $\hat{I}=\mathcal{F}(I,\hat{R}),\vspace{-0.2cm}$
where $\mathcal{F}(\cdot)$ denotes the deep network and $\hat{R}$ acts as an auxiliary guidance. Inspired by the improved Retinex model, we consider the enhancement process from two stages: Retinex-based guidance adjustment and conditional image enhancement. Recalling the conditional DDPM model discussed in Section 3.1, the enhancement task can be performed due to the superiority of the generative model conditioned by low-light images (Wang et al., 2023b; Jiang et al., 2023). However, the noise and artifacts in low-light images inevitably mislead the diffusion model and cause detail loss and unsatisfactory color.
@@
(8) $\hat{I}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(I,\mathcal{F_{A}}(R,L)),\mathcal{F_{A}}(R,L)),\vspace{-0.15cm}$
where $\bm{\epsilon}_{\theta}(\cdot)$, $\mathcal{F_{R}}(\cdot)$, and $\mathcal{F_{A}}(\cdot)$ denote UNet, RNet, and ANet in Fig. 3, respectively. Combining the promising texture generation capability of the diffusion model and the conscious representation with vivid color and detail of Retinex theory, we develop a Retinex-based condition strategy for the conditional DDPM and formulate our JoReS-Diff. Eq. 8 represents the denoising and refinement process in one iteration, which can be described as two stages, Retinex-based condition learning and Retinex-conditioned refinement. In the first stage, we obtain the decomposition results from a pre-trained DNet and input them into the following ANet to produce Retienx-based conditions, which will be described in Section 3.2.1. Then, the conditions will be introduced into the UNet and RNet to achieve conditional image denoising and refinement as elaborated in Section 3.2.2.
3.2.1. Retinex-based Condition Learning
As described in Section 3.2, the basis of conditional image refinement is to learn high-quality conditions. Inspired by Retinex-based methods (Wei et al., 2018; Zhang et al., 2019), we propose the DNet to serve as a pre-processing module and produce initial Retinex-based conditions. Compared to calculating an invariant color map as conditional input (Wang et al., 2023b; Yin et al., 2023), the decomposed maps contain not only color information but also details and illumination, achieving a more sufficient form of conditional inputs. To further exploit the progressively refined images in the diffusion process, DNet inputs $\mathbf{x}_{t}$ as well and provides competent updated predictions $R_{t}$ and $L_{t}$ for better generation. However, the images in the early steps contain incorrect contents and mislead the decomposition. Therefore, we use ANet to deal with corruption and ameliorate the low-quality conditions in the early steps, achieving approving conditional enhancement during the whole generative process. Thus, we realize the Retinex-based condition learning as:
(9) $\hat{\mathbf{c}}_{t}=\mathcal{F_{A}}(\mathbf{c}_{t})=\mathcal{F_{A}}(\mathcal{F_{D}}(\mathbf{y},\mathbf{x}_{t})).\vspace{-0.15cm}$
Pre-trained DNet. DNet adopts a lightweight UNet-like network to learn the decomposition mapping. It inputs the images $I$ and outputs the reflectance $R$ and illumination $L$. Inspired by the training strategy in (Wei et al., 2018), we utilize the constant reflectance loss, smooth illumination loss, and reconstruction loss to pre-train DNet.
@@ 3.2.1. Retinex-based Condition Learning
(10) $\mathcal{L}_{ja}=\parallel{R_{t}}\!^{\prime}-\hat{R}\parallel_{1}+SSIM({R_{t}}\!^{\prime},\hat{R}).\vspace{-0.1cm}$
@@
(11) $\mathcal{L}_{je}=\parallel{L_{t}}\!^{\prime}-\hat{L}\parallel_{1}+Hist({L_{t}}\!^{\prime},\hat{L}),\vspace{-0.15cm}$
@@
(12) $\mathcal{L}_{ANet}=\lambda_{ja}\mathcal{L}_{ja}+\lambda_{je}\mathcal{L}_{je},\vspace{-0.2cm}$
@@
3.2.2. Retinex-conditioned Refinement
After obtaining the Retinex-based conditions $\hat{\mathbf{c}}_{t}$, the next stage is to incorporate the conditions into the iterative diffusion process. Previous methods propose to introduce the invariant color maps (Yin et al., 2023; Wang et al., 2023b; Zhou et al., 2023) and Retinex decomposition (Yi et al., 2023). However, existing condition strategies are insufficient to conduct the favorable iterative enhancement. In our JoReS-Diff, we already learn the Retinex-based conditions as described in Section 3.2.1. Then, following the usage of color map (Yin et al., 2023; Wang et al., 2023b), we first control the generation capability of UNet by leveraging $\hat{\mathbf{c}}_{t}$ as extra inputs, while the carefully prepared conditions are being underutilized just for denoising. Thus, as illustrated in Eq. 8, we apply a post-processing refinement network (RNet) to fully exploit the conditions and guarantee the consistent color and contents of ${\hat{\mathbf{x}}_{0}}^{\prime}$. Different from (Yi et al., 2023), RNet adopts a lightweight architecture and implements Retinex-conditioned refinement at both feature- and image-levels by FRCM and IRCM. Overall, the refinement stage can be elaborated as:
(13) ${\hat{\mathbf{x}}_{0}}^{\prime}=\mathcal{F_{R}}(\bm{\epsilon}_{\theta}(\mathbf{x}_{t-1},\hat{\mathbf{c}}_{t}),\hat{\mathbf{c}}_{t}),\vspace{-0.2cm}$
@@ 3.2.2. Retinex-conditioned Refinement
Conditional Denoising. UNet integrates ${R_{t}}\!^{\prime}$ and ${L_{t}}\!^{\prime}$ by Retinex attention layers instead of input for effective interaction. We first transform ${R_{t}}\!^{\prime}$ and ${L_{t}}\!^{\prime}$ into embedding features $R_{em}$ and $L_{em}$. To maintain the content and color information, the reflectance embedding $R_{em}$ keeps the original resolution. The illumination embedding $L_{em}$ can be downsampled for computational efficiency thanks to the smooth distribution of lightness. Then, the $R_{em}$ and $L_{em}$ are input to Retinex attention layers with simple self-attention $F_{o}=SA(P_{em},F_{i}),\;P\!=\!R,L$, where $F_{i},F_{o}$ are input and output features. Thus, we can approximate $\hat{\mathbf{x}}_{0}$ by updating Eq. 2 as:
(14) $\hat{\mathbf{x}}_{0}=\frac{1}{\sqrt{\gamma_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\gamma_{t}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\hat{\mathbf{c}}_{t},\gamma_{t})\right).\vspace{-0.15cm}$
Figure 4. The architecture of the feature/image-level Retinex-conditioned modules (F/IRCM). FRCM inputs multi-scale features $F_{t}$ from ANet and obtains the optimized image feature $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$. Then, IRCM inputs the ${R_{t}}\!^{\prime}$, ${L_{t}}\!^{\prime}$ and the $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to refine the approximated $\hat{\mathbf{x}}_{0}$ a produce the final output ${\hat{\mathbf{x}}_{0}}^{\prime}$.
Conditional Refinement. As shown in Fig. 3, RNet consists of two Retinex-conditioned modules to realize the color recovery and detail enhancement of the approximated $\hat{\mathbf{x}}_{0}$. Specifically, we first implement the Retinex-conditioned refinement in feature space, since $F_{t}$ incorporates rich information of adjustment mapping as illustrated in Section 3.2.1. Thus, as depicted in Fig. 4, we unify the sizes of multi-scale features in $F_{t}$ and model them as affine transformation parameters $F_{\gamma}$ and $F_{\beta}$. Then, the feature fusion is carried out by scaling and shifting operations $\mathcal{F}_{T}(\cdot|\cdot)$ as:
(15) $F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}=\mathcal{F}_{T}(F_{\hat{\mathbf{x}}_{0}}|F_{t})=F_{\gamma}\odot W(F_{\hat{\mathbf{x}}_{0}})+F_{\beta},\vspace{-0.15cm}$
@@
(16) $\begin{split}\Delta\hat{\mathbf{x}}_{0}&=\mathcal{F}_{T}({R_{t}}\!^{\prime}|\mathcal{F}_{T}({L_{t}}\!^{\prime}|\hat{\mathbf{x}}_{0})),\\ {\hat{\mathbf{x}}_{0}}^{\prime}&=\mathcal{F}_{T}(\hat{\mathbf{x}}_{0}|\Delta\hat{\mathbf{x}}_{0})+W(F_{\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}),\vspace{-0.3cm}\end{split}$
@@
where ${\hat{\mathbf{x}}_{0}}^{\prime}$ is the final refined output of RNet. In RNet, FRCM and IRCM fully exploit the Retinex-based conditions and produce results with consistent color and details. Accordingly, in addition to Eq. 4, we also utilize the constraint of ${\hat{\mathbf{x}}_{0}}^{\>{}^{\prime}}$ to optimize the JoReS-Diff as:
@@
(17) $\mathcal{L}_{RNet}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\mathbf{x}_{0}-\mathcal{F}_{R}(\hat{\mathbf{x}}_{0},\hat{\mathbf{c}}_{t},\gamma_{t})\rVert^{2}_{2}].\vspace{-0.15cm}$
3.3. Semantic Prior Incorporation
Apart from the capability of color recovery and detail maintenance brought by the Retinex prior, we introduce the semantic prior to further control the generative diffusion model at semantic-level and alleviate the unnatural textures in the output image. Following the semantic-aware framework in (Wu et al., 2023), we utilize the same semantic segmentation model to produce available semantic prior from the input image. The segmentation model inputs the low-light image and outputs the intermediate features and semantic map. However, the low-light image remains the distribution gap comparing to the normal-light image, which may cause the misclassification and thus produce unsatisfactory segmentation results. Directly applying the semantic map will mislead the diffusion model in some cases, resulting in more unnatural textures. Therefore, we only use the intermediate latent features since the misclassification can be smoothed in the high-dimensional feature space. Consequently, we collect the multi-scale features from the segmentation model as the semantic prior, $\mathbf{c}_{seg}=[{F_{s}}^{0},{F_{s}}^{1},{F_{s}}^{2}]$. The three semantic features (${F}_{s}^{b},b=0,1,2$) with three spatial resolutions ($\textit{H}/2^{4-b},\textit{W}/2^{4-b}$), where H and W are the height and width of the input image.
Then, we propose the semantic attention layers to conduct the introduction of the semantic prior, as shown in Fig. 3. Similar to the Conditional Denoising in Section 3.2.2, we exploit the semantic information to manipulate the feature in the decoder of the UNet as well. Notably, we design the semantic attention layer based on both the self-attention $SA(\cdot)$ and cross-attention $CA(\cdot)$. Primarily, self-attention plays a crucial role in preserving the geometric and shape details and the cross-attention contributes more to generate semantic consistency (Liu et al., 2024). And we apply learnable weights to achieve an adaptive addition of the outputs from self- and cross-attention branches. The overall calculation can be described as:
(18) $F_{o}=\lambda_{SA}SA(F_{s},F_{i})+\lambda_{CA}CA(F_{s},F_{i}),\vspace{-0.25cm}$
where $\lambda_{SA},\lambda_{CA}$ denote learnable weights, $F_{i},F_{o}$ denote input and output features. Consequently,  Eq. 4 can be reformulated as:
(19) $\begin{split}\mathcal{L}_{UNet}&=\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},t,\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\lVert\bm{\epsilon}_{t}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},\mathbf{y},\hat{\mathbf{c}}_{t},\mathbf{c}_{seg},\gamma_{t})\rVert^{2}_{2}],\\ \mathcal{L}_{ALL}&=\lambda_{UNet}\mathcal{L}_{UNet}+\lambda_{RNet}\mathcal{L}_{RNet}+\lambda_{ANet}\mathcal{L}_{ANet}\vspace{-0.25cm}\end{split}$
where $\lambda_{s}$ denote loss weights.
4. Experiments
4.1. Experimental Settings
@@
Table 1. Quantitative comparison on the LOL (Wei et al., 2018) and LOL-v2-real (Yang et al., 2021) datasets. $\uparrow$ ($\downarrow$) denotes that, larger (smaller) values suggest better quality. The best results are highlighted in bold and the second best results are in underline (Special fonts are also used in Table 2). Note that the absent results (“-”) of Diff-Retinex due to the lack of code.
Methods LOL LOL-v2-real LOL-v2-synthetic
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
LIME (Guo et al., 2016) TIP’16 16.760 0.560 0.350 117.892 15.240 0.470 0.428 118.171 16.880 0.758 0.104 -
RetinexNet (Wei et al., 2018) BMVC’18 16.770 0.462 0.474 113.699 18.371 0.723 0.365 133.905 16.551 0.652 0.379 98.843
KinD (Zhang et al., 2019) MM’19 20.870 0.799 0.207 104.632 17.544 0.669 0.375 137.346 18.956 0.801 0.262 89.156
DRBN (Yang et al., 2020) CVPR’20 19.860 0.834 0.155 75.359 20.130 0.830 0.147 60.631 21.687 0.825 0.174 52.972
Zero-DCE (Guo et al., 2020) CVPR’20 14.861 0.562 0.335 101.237 18.059 0.580 0.313 91.939 17.756 0.814 0.168 49.239
MIRNet (Zamir et al., 2022b) PAMI’22 24.140 0.842 0.131 69.179 20.357 0.782 0.317 49.108 21.941 0.876 0.112 38.775
SNR (Xu et al., 2022) CVPR’22 24.608 0.840 0.151 55.121 21.479 0.848 0.157 54.532 24.130 0.927 0.032 23.971
PairLIE (Fu et al., 2023) CVPR’23 19.510 0.736 0.248 100.715 20.357 0.782 0.317 96.911 19.074 0.794 0.230 85.209
SMG (Xu et al., 2023) CVPR’23 23.684 0.826 0.118 58.846 24.620 0.867 0.148 78.582 25.618 0.905 0.053 23.210
FourLLIE (Li et al., 2023a) MM’23 20.222 0.766 0.250 91.793 22.340 0.847 0.051 89.334 24.649 0.919 0.039 26.351
Retinexformer (Cai et al., 2023) ICCV’23 25.153 0.843 0.131 71.148 22.794 0.839 0.171 62.439 25.670 0.928 0.059 22.781
Diff-Retinex (Yi et al., 2023) ICCV’23 21.981 0.863 0.048 47.851 - - - - - - - -
DiffLL (Jiang et al., 2023) SIGGRAPH Asia’23 26.336 0.845 0.217 48.114 22.428 0.817 0.191 59.075 25.456 0.896 0.102 43.670
PyDiff (Zhou et al., 2023) IJCAI’23 25.643 0.849 0.142 69.784 23.441 0.833 0.208 71.538 25.126 0.917 0.098 29.361
JoReS-Diff (Ours) 26.491 0.876 0.092 43.596 24.836 0.897 0.109 46.938 25.712 0.928 0.058 22.776
@@
Table 2. Quantitative comparison on the UHD-LL (Li et al., 2023a). To conserve space, we select recent methods in 2023 while ensuring that the results are sufficient to demonstrate our superiority.